# SwitchLoRA: Switched Low-Rank Adaptation Can Learn Full-Rank Information

### v2

@@ 4.1 Experimental setup
Our studies are carried out on the LLaMA model [45], with model sizes reduced to 130M, 250M, and 350M. We designed our experiments based on the settings described in [34] to benefit from established hyperparameter configurations. The specific hyperparameters for these models are detailed in Table 1. We use Adam optimizer to train the model with $\beta_{1}=0.9,\beta_{2}=0.999$. We use a cosine learning rate schedule with 100 warm-up steps and a total of 40,000 training steps.
The pre-training experiments utilize the C4 dataset [9], with the first 46M samples of the training dataset serving as our training data, and samples from the entire validation dataset used for testing. The evaluation of validation loss is performed on 10M tokens for all our experiments, with evaluations conducted every 1,000 steps. Additionally, we utilize some of tasks from the GLUE benchmark [48] to assess the reasoning capabilities of the models. All experiments are conducted using 8xNVIDIA Tesla A100 80GB PCIe GPUs. Gradient accumulation is applied when GPU memory reaches its limit.
@@
Table 1: Model sizes and architectures used in our experiments
@@ 4.2 Basic experiments
Table 2: Perplexity results at step 38,000 for 130M, 250M and 350M.
@@
Table 3: Perplexity results at step 38,000 for 1.3B models.
@@ 4.3 Comparison with other methods
Among all related methods, the works which are most close to ours are ReLoRA [34] and GaLore [60]. We do comparison experiments on these two methods to further validate the effectiveness of our algorithm. The learning rates for all methods are tuned in $\cup_{n=2,3,4}${1e-n 2e-n, 5e-n}.
@@
Figure 4: Comparison between ReLoRA and SwitchLoRA. In the figure, red circles denotes the steps at which the parameters of the LoRA adapter are reset.
@@
In the comparison experiments with GaLore, we strictly follow the setup in Galore [60]. Detailed setup can be found in Appendix C. For the 350M LLaMA model, GaLore achieves a perplexity of 20.29, whereas SwitchLoRA performs slightly better, with a perplexity of 19.58. In addition, we conducted additional experiments on the 350M model, changing only one hyperparameter to assess its impact. The perplexity results are shown in Table 5.
When further reducing the rank, as shown in Table 5, our method performs significantly better. This improvement may be because GaLore’s use of SVD focuses on the most significant directions, whereas SwitchLoRA covers all update directions, including less important ones that still require training.
Table 5: Perplexity comparison for GaLore and SwitchLoRA with different experimental setup.
@@ 4.4 Reasoning ability comparison
Current works on low-rank training for LLMs, such as [34, 60], primarily evaluate models based on perplexity and lack validation of reasoning abilities. To validate the reasoning abilities, we also conducted full fine-tuning using the resulting checkpoints from the aforementioned experiments. We fine-tuned the models on GLUE tasks [48]. For the checkpoints trained using SwitchLoRA, all LoRA adapters are merged into the original weights such that $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{B}\mathbf{A}$ before the fine-tuning process. Detailed experiment settings are provided in Appendix C.
@@
CoLA STS-B MRPC RTE SST2
Full-rank pre-trained 42.95$\pm$5 87.26$\pm$0.2 79.16$\pm$1 59.86$\pm$1 90.88$\pm$0.5
SwitchLoRA pre-trained 23.13$\pm$15 87.71$\pm$0.5 76.86$\pm$2 56.24$\pm$5 90.83$\pm$0.3
GaLore pre-trained 40.23$\pm$2 86.14$\pm$0.5 72.70$\pm$4 54.66$\pm$4 89.35$\pm$0.5
@@
The full fine-tuning results for these three models are shown in Table 6. From the results, we observe that for the 350M models, except for the CoLA task, SwitchLoRA outperforms GaLore by [insert performance difference], and outperforms the full-rank model by [insert performance difference].
We also conduct fine-tuning experiments on the 1.3B LLaMA models, one pre-trained using full-rank and the other pre-trained using SwitchLoRA with a LoRA rank of 512. Both pre-trained models are from Section 4.2. As shown in Table 7, SwitchLoRA performs slightly worse in some tasks and better in others compared to the full-rank results. Overall, the average score of SwitchLoRA exceeds the full-rank results by approximately 1%.
Table 7: GLUE benchmark of the Full-rank pre-trained and SwitchLoRA pre-trained models. The metrics for all tasks except STS-B are accuracy. The metric for STS-B is Pearson correlation coefficient.
@@ 5 Limitations and future work
  • Going further, a more detailed idea is to examine each layer of the model to adjust the switching frequencies. For instance, LoRA-drop [61] evaluates whether the rank is sufficient using a norm of $\Delta\mathbf{W}\mathbf{x}$. This is rational because different types of layers, such as the $Q,K,V$ matrices in transformer layers, exhibit significantly varied behaviors.
@@ 6 Conclusions
In this work, we introduce SwitchLoRA, a novel training strategy designed for parameter-efficient pre-training. Our approach achieves comparable accuracy to full-rank training while reducing the trainable parameters to approximately 50% to 60% of those in traditional full-rank training. Moreover, the computational overhead and memory usage are nearly identical to those of LoRA when using the same number of trainable parameters. We further validate the reasoning abilities of models trained with SwitchLoRA using the GLUE benchmark. The results from the 1.3B model indicate that SwitchLoRA not only matches but also slightly outperforms full-rank training by about 1% in accuracy.
@@ Appendix A Theoretical analysis
In optimization, it is well-established that for problems with separable objective functions, the parameters of each separable group can be optimized independently. Although the loss function of the SwitchLoRA model is not separable, the preceding discussion has demonstrated the independence between the LoRA vectors. Consequently, we can infer that the inseparable components of the loss function concerning parameters within the same linear layer are modest. Therefore, this suggests that training subsets of parameters incrementally, as in the SwitchLoRA model, is likely more effective than other methods, such as the layer-wise training approach [3, 2].
@@
The main idea of [12] and [17] is to maintain a balance in the variance of the activation and gradients across layers during forward and backward propagation. In this study, we focus on balancing the variance of activations. Furthermore, we aim to ensure the updated parameters derived from $\mathbf{B}$ are of the same amount as those derived from $\mathbf{A}$:
@@ Appendix C Experimental setting details
C.1 Experimental settings of ReLoRA
We adhere strictly to the setup described in ReLoRA [34] for our comparative experiments with ReLoRA. Specifically, the warm-up steps for the scheduler are set to 1,000. The learning rates are as follows: 5e-4 for full-rank pretraining, 1e-3 for ReLoRA, and 1e-2 for SwitchLoRA. The total batch size is established at 20,000. All other settings remain consistent with our previous experiments as detailed in Section 4.
C.2 Experimental settings of GaLore
Continuing in the same vein, we also strictly adhere to the setup outlined in GaLore [60] for our comparison experiments with GaLore. Specifically, we set the warm-up steps for the scheduler at 6,000. The total batch size is adjusted to 60,000. The learning rate is standardized at 1e-2 for all GaLore experiments, while for SwitchLoRA, it is set at 2e-2. All other experimental settings remain consistent with those detailed in our previous experiments, as described in Section 4.
C.3 Experimental settings of fine-tuning
@@
CoLA STS-B MRPC RTE SST2
$lr$ (Full-rank) 8e-6 1e-5 1e-5 8e-6 3e-6
$lr$ (SwitchLoRA) 3e-5 5e-5 5e-5 5e-5 1e-5
$lr$ (GaLore) 2e-6 5e-6 5e-6 3e-6 8e-7
@@ Appendix D Implementation of LoRA vector switching
The primary distinction in the implementation of SwitchLoRA from conventional approaches lies in its handling of gradients and optimizer states at the granularity of row or column vectors within matrix parameters. Consider the scenario when using the AdamW optimizer: typically, each trainable parameter group in AdamW is associated with a “step” state which is implemented as a float scalar value in the code. To facilitate the resetting of specific rows or columns in matrices, we modify the type of "step" in the optimizer to a 32-bit float matrix with the same shape as the corresponding parameters. In fact, this modification does incur some extra memory overhead. An alternative approach would be to implement “step” as a row vector for $\mathbf{A}$ and a column vector for $\mathbf{B}$. However, this would require more complex code management, and thus, we have not adopted this strategy in our implementation. With the capability to manipulate optimizer states and gradients at the level of rows and columns, we can now execute operations such as resetting optimizer states and freezing specific rows or columns of parameter matrices.
@@ Appendix E Distribution of Singular values
Given that the rank distribution significantly influences the training efficacy of models [21, 11], we conducted experiments to examine the rank distribution of SwitchLoRA. As outlined in Section 4, experiments were conducted on the 350M model to analyze the rank distribution of linear layers after 40,000 training steps. Figure 10 demonstrates that the singular values of weight matrices converge within a limited range when trained with LoRA, indicating dominance of LoRA adapters in the linear layers. This dominance is expected, as the singular value distribution of weight matrices during the pre-training phase exhibits a form of illness, due to updates being limited to the low-rank adapter $\mathbf{B}\mathbf{A}$. In contrast, as illustrated in Figure 11, the rank distribution of SwitchLoRA closely approximates that of full-rank training, suggesting a more robust and more effective adaptation process.
@@ Appendix F Impact on distributed training
As demonstrated in [40], for a transformer model with $n$ layers and a hidden dimension of $h$, the memory required for model parameters scales proportionally with $nh^{2}$. Assuming these parameters are stored in $fp16/bf16$ format occupying $\Psi$ parameters, the memory footprint for optimizer states would be approximately $12\Psi$ bytes when using the Adam optimizer as stated in [40]. Additionally, when the batch size is $b$ and the sequence length is $s$, the memory consumption for activations scales with $bshn$. To manage memory demands for large models, gradient accumulation can be utilized to adjust the batch size per GPU to $1$. Moreover, activation checkpointing can be implemented to reduce memory consumption, though it comes with a trade-off: a 33% increase in computational overhead.

### v3

@@ 4.1 Experimental setup
Our studies are carried out on the LLaMA model (Touvron et al., 2023), with model sizes reduced to 130M, 250M, and 350M. We designed our experiments based on the settings described in Lialin et al. (2023) to benefit from established hyperparameter configurations. The specific hyperparameters for these models are detailed in Table 1. We use Adam optimizer to train the model with $\beta_{1}=0.9,\beta_{2}=0.999$. We use a cosine learning rate schedule with 100 warm-up steps and a total of 40,000 training steps.
The pre-training experiments utilize the C4 dataset (Dodge et al., 2021), with the first 46M samples of the training dataset serving as our training data, and samples from the entire validation dataset used for testing. The evaluation of validation loss is performed on 10M tokens for all our experiments, with evaluations conducted every 1,000 steps. Additionally, we utilize some of tasks from the GLUE benchmark (Wang et al., 2019) to assess the reasoning capabilities of the models. The experiments are conducted using 8xNVIDIA A800 80GB PCIe GPUs. Gradient accumulation is applied when GPU memory reaches its limit.
@@
Table 1: Model sizes and architectures used in our experiments.
@@ 4.2 Basic experiments
Table 2: Perplexity results for 130M, 250M and 350M.
@@
Table 3: Perplexity results for 1.3B models.
@@
Memory usage and training time
In larger models, the communication overhead from data parallelism and memory usage for optimizer states become the dominant factors, as the batch size per GPU decreases. Table 5 compares memory usage and training time for the 1.3B, 3B, and 7B LLaMA models. Detailed hyperparameters for the 3B and 7B models are provided in Appendix C.
Table 5: Memory usage and training comparison of full-rank training and LoRA/SwitchLoRA training on 4 A800 PCIe GPUs. The total batch size is 1536. “bs” refers to the number of samples processed in each substep before accumulating gradients. We record the training time and the amount of memory offloaded for candidate vectors to the CPU for one step. The LoRA rank for both LoRA and SwitchLoRA is set to $hidden\_dim/4$, and the switching frequency of SwitchLoRA is set to 1/40.
Model Size Method Time (sec) Trainable Param. Memory Usage Offlaoded Memory
[…]
SwitchLoRA 22.5 610M 31.9GB 95.6MB
3B (bs=4) Full-rank 48.4 2686M 37.4GB
[…]
SwitchLoRA 39.4 1162M 27.1GB 199MB
[…]
SwitchLoRA 194 2822M 47.3GB 512MB
@@ 4.3 Comparison with other methods
Among all related methods, the works which are most close to ours are ReLoRA (Lialin et al., 2023) and GaLore (Zhao et al., 2024b). We do comparison experiments on these two methods to further validate the effectiveness of our algorithm. The learning rates for all methods are tuned in $\cup_{n=2,3,4}${1e-n 2e-n, 5e-n}.
@@
Figure 4: Comparison between ReLoRA and SwitchLoRA. In the figure, red circles denotes the steps at which the parameters of the LoRA adapter are reset. In the left figure, ReLoRA utilizes 5,000 steps of full-rank pre-training, while SwitchLoRA uses 200 steps. In the right figure, both algorithms employ 1,000 steps of full-rank pre-training.
@@
We also test ReLoRA with a higher resetting frequency during the early training phase, similar to SwitchLoRA. However, the training loss tends to either diverge or decrease slowly.
@@
In the comparison experiments with GaLore, we strictly follow the setup in Galore (Zhao et al., 2024b). Detailed setup can be found in Appendix C. For the 350M LLaMA model, GaLore achieves a perplexity of 20.29, whereas SwitchLoRA performs slightly better, with a perplexity of 19.58. In addition, we conducted additional experiments on the 350M model, changing only one hyperparameter to assess its impact. The perplexity results are shown in Table 6.
When further reducing the rank, as shown in Table 6, our method performs significantly better. This improvement may be because GaLore’s use of SVD focuses on the most significant directions. As a result, there is an inherent loss of information during gradient compression. In contrast, SwitchLoRA covers all update directions, including less important ones that still require training.
Table 6: Perplexity comparison for GaLore and SwitchLoRA with different experimental setup. The standard result is from the 350M model with a rank of 256 and a sequence length of 256.
@@ 4.4 Reasoning ability comparison
Current works on low-rank training for LLMs, such as Lialin et al. (2023); Zhao et al. (2024b), primarily evaluate models based on perplexity and lack validation of reasoning abilities. To validate the reasoning abilities, we also conducted full fine-tuning using the resulting checkpoints from the aforementioned experiments. We fine-tuned the models on GLUE tasks (Wang et al., 2019). For the checkpoints trained using SwitchLoRA, all LoRA adapters are merged into the original weights such that $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{B}\mathbf{A}$ before the fine-tuning process. Detailed experiment settings are provided in Appendix C.
@@
CoLA STS-B MRPC RTE SST2 MNLI QNLI QQP
Full-rank pre-trained 43.0$\pm$5 87.3$\pm$0.2 79.2$\pm$1 59.9$\pm$1 90.9$\pm$0.5 80.1$\pm$0.3 85.30$\pm$0.1 89.8 $\pm$0.04
SwitchLoRA pre-trained 43.3$\pm$3 87.8$\pm$0.3 77.0$\pm$2 61.9$\pm$4 90.1$\pm$1 81.7$\pm$0.3 86.6$\pm$1 89.7$\pm$0.2
GaLore pre-trained 40.2$\pm$2 86.1$\pm$0.5 72.7$\pm$4 54.7$\pm$4 89.4$\pm$0.5 77.8$\pm$0.4 84.6$\pm$0.4 89.0$\pm$0.1
@@
The full fine-tuning results for these three models are shown in Table 7. From the results, we observe that for the 350M models, SwitchLoRA outperforms GaLore by an average score of around 3.0 points, and outperforms the full-rank model by an average score of around 0.3 points.
We also conduct fine-tuning experiments on the 1.3B LLaMA models, one pre-trained using full-rank and the other pre-trained using SwitchLoRA with a LoRA rank of 512. Both pre-trained models are from Section 4.2. As shown in Table 8, SwitchLoRA performs slightly worse in some tasks and better in others compared to the full-rank results. Overall, the average score of SwitchLoRA exceeds the full-rank results by approximately 1.0 points.
Table 8: GLUE benchmark of the full-rank and SwitchLoRA pre-trained 1.3B models. The metric for STS-B is the Pearson correlation, while Matthew’s correlation coefficient is used for CoLA. Accuracy is reported for the other tasks.
@@ 5 Limitations and future work
  • Going further, a more detailed idea is to examine each layer of the model to adjust the switching frequencies. For instance, LoRA-drop (Zhou et al., 2024) evaluates whether the rank is sufficient using a norm of $\Delta\mathbf{W}\mathbf{x}$. This is rational because different types of layers, such as the $Q,K,V$ matrices in transformer layers, exhibit significantly varied behaviors.
@@ 6 Conclusions
In this work, we introduce SwitchLoRA, a novel training strategy designed for parameter-efficient pre-training. Our approach achieves comparable accuracy to full-rank training while reducing the trainable parameters to approximately 50% to 60% of those in traditional full-rank training, significantly decreasing communication time in data parallelism. Moreover, the computational overhead and memory usage are nearly identical to those of LoRA when using the same number of trainable parameters. We further validate the reasoning abilities of models trained with SwitchLoRA using the GLUE benchmark. The results from the 1.3B model indicate that SwitchLoRA not only matches but also slightly outperforms full-rank training by about 1% in accuracy.
@@ Appendix A Theoretical analysis
In optimization, it is well-established that for problems with separable objective functions, the parameters of each separable group can be optimized independently. Although the loss function of the SwitchLoRA model is not separable, the preceding discussion has demonstrated the independence between the LoRA vectors. Consequently, we can infer that the inseparable components of the loss function concerning parameters within the same linear layer are modest. Therefore, this suggests that training subsets of parameters incrementally, as in the SwitchLoRA model, is likely more effective than other methods, such as the layer-wise training approach (Bengio et al., 2006; Allen-Zhu and Li, 2020).
@@
The main idea of Glorot and Bengio (2010) and He et al. (2015) is to maintain a balance in the variance of the activation and gradients across layers during forward and backward propagation. In this study, we focus on balancing the variance of activations. Furthermore, we aim to ensure the updated parameters derived from $\mathbf{B}$ are of the same amount as those derived from $\mathbf{A}$:
@@ Appendix C Experimental setting details
C.1 Model sizes and architectures of larger models
Table 9 presents the hyperparameters of the model architectures used to evaluate training time and memory usage, as reported in Table 5.
Table 9: Model sizes and architectures used to evaluate training time and memory usage.
Params Hidden Heads Layers Batch size Batch size per GPU Seq. len.
1.3B 2048 32 24 1536 16 512
3B 2560 32 32 1536 4 512
7B 4096 32 32 1536 1 512
C.2 Experimental settings of ReLoRA
We adhere strictly to the setup described in ReLoRA (Lialin et al., 2023) for our comparative experiments with ReLoRA. Specifically, the warm-up steps for the scheduler are set to 1,000. The learning rates are as follows: 5e-4 for full-rank pre-training, 1e-3 for ReLoRA, and 1e-2 for SwitchLoRA. The total batch size is established at 20,000. All other settings remain consistent with our previous experiments as detailed in Section 4.
C.3 Experimental settings of GaLore
Continuing in the same vein, we also strictly adhere to the setup outlined in GaLore (Zhao et al., 2024b) for our comparison experiments with GaLore. Specifically, we set the warm-up steps for the scheduler at 6,000. The total batch size is adjusted to 60,000. The learning rate is standardized at 1e-2 for all GaLore experiments, while for SwitchLoRA, it is set at 2e-2. All other experimental settings remain consistent with those detailed in our previous experiments, as described in Section 4.
C.4 Experimental settings of fine-tuning
@@
CoLA STS-B MRPC RTE SST2 MNLI QNLI QQP
$lr$ (Full-rank) 8e-6 1e-5 1e-5 8e-6 3e-6 2e-5 1e-5 2e-5
$lr$ (SwitchLoRA) 1e-5 3e-5 3e-5 1e-5 8e-6 3e-5 3e-5 3e-5
$lr$ (GaLore) 2e-6 5e-6 5e-6 3e-6 8e-7 8e-6 3e-6 5e-6
Epochs 30 30 30 30 30 10 10 10
@@ Appendix D Implementation of LoRA vector switching
The primary distinction in the implementation of SwitchLoRA from conventional approaches lies in its handling of gradients and optimizer states at the granularity of row or column vectors within matrix parameters. Consider the scenario when using the AdamW optimizer: typically, each trainable parameter group in AdamW is associated with a “step” state which is implemented as a float scalar value in the code. To facilitate the resetting of specific rows or columns in matrices, we modify the type of “step” in the optimizer states to a row vector for LoRA matrices $\mathbf{A}$ and a column vector for LoRA matrices $\mathbf{B}$.
To be more specific, consider the parameters “param” and their corresponding optimizer states “state”,. The original PyTorch code to initialize the optimizer state “step” is as follows:
{python}
state["step"] = torch.zeros(1, 1, dtype=torch.float32)
The modified code is:
{python}
if is_lora_A(param): state["step"] = torch.zeros(param.shape[0], 1, dtype=torch.float32) elif is_lora_B(param): state["step"] = torch.zeros(1, param.shape[1], dtype=torch.float32) else: state["step"] = torch.zeros(1, 1, dtype=torch.float32)
With the capability to manipulate optimizer states and gradients at the level of rows and columns, we can now execute operations such as resetting optimizer states and freezing specific rows or columns of parameter matrices.
@@ Appendix E Distribution of Singular values
Given that the rank distribution significantly influences the training efficacy of models (Hu et al., 2022; Frankle and Carbin, 2019), we conducted experiments to examine the rank distribution of SwitchLoRA. As outlined in Section 4, experiments were conducted on the 350M model to analyze the rank distribution of linear layers after 40,000 training steps. Figure 10 demonstrates that the singular values of weight matrices converge within a limited range when trained with LoRA, indicating dominance of LoRA adapters in the linear layers. This dominance is expected, as the singular value distribution of weight matrices during the pre-training phase exhibits a form of illness, due to updates being limited to the low-rank adapter $\mathbf{B}\mathbf{A}$. In contrast, as illustrated in Figure 11, the rank distribution of SwitchLoRA closely approximates that of full-rank training, suggesting a more robust and more effective adaptation process.
@@ Appendix F Impact on distributed training
As demonstrated in Rajbhandari et al. (2020), for a transformer model with $n$ layers and a hidden dimension of $h$, the memory required for model parameters scales proportionally with $nh^{2}$. Assuming these parameters are stored in $fp16/bf16$ format occupying $\Psi$ parameters, the memory footprint for optimizer states would be approximately $12\Psi$ bytes when using the Adam optimizer as stated in Rajbhandari et al. (2020). Additionally, when the batch size is $b$ and the sequence length is $s$, the memory consumption for activations scales with $bshn$. To manage memory demands for large models, gradient accumulation can be utilized to adjust the batch size per GPU to $1$. Moreover, activation checkpointing can be implemented to reduce memory consumption, though it comes with a trade-off: a 33% increase in computational overhead.